# Identifying treatment response subgroups in observational time-to-event data

### v1

@@ 3.2 Empirical settings
Benchmarks.
We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE [34] [^7] Implemented in the Auton-Survival library [35]) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach is that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies subgroups of treatment while considering survival, without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, in comparison to a group with the same treatment responses but better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation.
Finally, we also compare our model against its unadjusted alternative (CNSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
Training.
We perform a 5-fold cross-validation for both Randomised and Observational simulations. For each cross-validation split, the development set is divided into three parts: 80% for training, 10% for early stopping, and 10% for hyper-parameter search over a grid presented in Appendix B.2. All models were optimised for 1000 epochs using an Adam optimiser [23] with early stopping.
@@
As, in the synthetic experiments, the subgroup structure is known, we measure the adjusted [^8] Random patient assignment results in an adjusted Rand-Index of 0. Rand-Index [39], which quantifies how the estimated assignment aligns with the known underlying group structure. Additionally, we use the integrated squared error (ISE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
@@
where $\hat{\tau}_{k}$ is the estimated treatment effect for subgroup $k$, and $\tau_{k}$ is the ground truth. Our experiments use a numerical approximation of the integral at the last observed event time. Further, we measure the ISE at the population level, defined as
[…]
@@ 3.3 Treatment effect recovery
A limitation of our approach is that we need to choose the number of subgroups a priori. An important question to analyse is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic [50], which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, aligning with the underlying generative process. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses, which first model the treatment effect and subsequently cluster the population. Consequently, survival outcomes do not directly guide the choice of $K$ in classical approaches.
@@
Next, we present quantitative results of our method against benchmarks. Table 1 presents the performance of the different methodologies on the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
[…]
@@
Table 1: Cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. ISEpop quantifies the error in estimating the average treatment effect, while ISEk reflects how the estimated treatment effects align with the ground truth. Best performance per column and simulation scenario are marked in bold. Our proposed CNSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
@@
Our proposed CNSC methods outperform all alternatives of CMHE, the current state-of-the-art method, in the proposed experiments. We hypothesise CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards. Neither of these two assumptions is likely to hold in real-world settings. In contrast, our approach does not constrain the treatment effect through flexible modelling of the survival function under both treatment regimes.
Increasing the number of subgroups, as shown in $K=L$, improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower ISEpop) but still underperforms in comparison to our proposed methods. These experiments highlight the advantage of CNSC in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
Moreover, the Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. While all methodologies present larger ISE in the observational setting, we observe a larger performance gap between CMHE and CNSC, highlighting the importance of accounting for the treatment assignment bias. Further, the two CNSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. The superiority of CNSC over its unadjusted alternative is evident in the observational setting, where CNSC’s average treatment effect estimation performance is better, particularly in the recovery of treatment effect of subgroup 3, where not adjusting for treatment assignment doubles the ISE.
@@ 4 Related work
While survival subgrouping has been proposed through mixtures of distributions [32, 33, 19] to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current literature focuses on estimating population or individual treatment effects [9, 45, 22, 55]. While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
@@
Discovering intervenable subgroups is core to medical practice, particularly identifying subgroups of treatment effect, as patients do not respond like the average [6, 43, 44] and the average may conceal differential treatment responses. Identification of subgroups has long been used to design RCTs. Indeed, subgroups identified a priori can then be tested through trials [8, 41]. A posteriori analyses have gained traction to uncover subgroups of patients from existing RCTs to understand the underlying variability of responses.
@@
The first set of subgroup analysis methodologies consists of a step-wise approach: (i) estimate the ITE and (ii) uncover subgroups using a second model to explain the heterogeneity in ITE. Foster et al. [11], Qi et al. [38] describe the virtual twins approach in which one models the outcome using a decision tree for each treatment group. The difference between these decision trees results in the estimated treatment effects. A final decision tree aims to explain these estimated treatment effects to uncover subgroups. Similar approaches have been explored with different meta-learners [53], or Bayesian additive models [17], or replacing the final step with a linear predictor to uncover the feature influencing heterogeneity [7]. However, Guelman et al. [12] discuss the drawbacks associated with these approaches. Notably, the two-step optimisation may not lead to recovery of the underlying subgroups of treatment effects.
@@
Tree-based approaches have been proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. [49] introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey and Imbens [2, 1] improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager and Athey [51] agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect [27, 28]. Alternatively, McFowland III et al. [30] propose pattern detection and Wang and Rudin [52], causal rule set learning to uncover these subgroups. However, all these approaches rely on a criticised local optimisation criterion [26] and greedy split exploration. Recently, Nagpal et al. [31] addressed the local optimisation by constraining the treatment response shape to a linear form in a mixture of Cox models.
@@
The previous approaches uncover subgroups of treatment effect but often consider RCTs with binary outcomes, not the observational setting with the inherently continuous survival outcomes that we are interested in. At the intersection with survival analysis, Zhang et al. [54] extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. [17] propose Bayesian additive models and Zhu and Gallego [56] propose a step-wise approach with propensity weighting to study observational data. Closest to our work, [21, 36] propose to uncover subgroups with survival outcomes. Specifically, Nagpal et al. [36] stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As discussed, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data. Similarly, Jia et al. [21] analyse RCTs with a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework.
@@ 5 Conclusion
This work introduces a novel methodology to uncover subgroups of treatment responses in observational time-to-event data. Our methodology addresses, to date, unaddressed challenges in subgroup discovery, specifically non-randomised treatment assignment and censoring. While modelling the observed outcome, the approach simultaneously identifies patient subgroups with different treatment responses in a joint approach. These groups may motivate the design of future clinical trials and support the development of guidelines to better target treatment to subgroups that would most benefit from them.
Our experiments demonstrate the capacity of our proposed method to uncover underlying groups of treatment effects. However, causal modelling relies on empirically unverifiable assumptions. The invalidity of these assumptions is a lesser concern in hypothesis generation, as offered by the proposed model, than in individualised estimates aimed at treatment recommendations.
Serving the purpose of a hypothesis-generating tool, we invite practitioners to further study the subgroups, identified in observational studies, in RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
Acknowledgments and Disclosure of Funding
VJ acknowledges the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2) and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1. FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z), and the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1.

### v3

@@ 3.2 Empirical settings
Benchmark methods.
We compare the proposed approach against the state-of-the-art Cox Mixtures with Heterogeneous Effect (CMHE Nagpal et al. (2022a) [^5] Implemented in the Auton-Survival library (Nagpal et al., 2022b)) which uncovers treatment effect and baseline survival latent groups. This method uses an expectation maximisation framework in which each patient is assigned to a group for which a Cox model is then fitted. The central differences with our proposed approach is that CMHE (i) clusters patients in treatment effects subgroups and survival subgroups, and (ii) assumes a linear treatment response. This separation between survival and treatment response improves the model’s flexibility at the cost of interpretability as the number of groups grows exponentially, and subgroups of treatment effect are independent of survival. Further, the assumption of linear treatment response may hinder the discovery of subgroups with more complex responses. By contrast, our approach identifies subgroups of treatment while considering survival, without constraining the treatment response. We argue that these are key strengths of our method as a group not responding to treatment with low life expectancy would most benefit from alternative treatments, in comparison to a group with the same treatment responses but better survival odds. Considering jointly non-linear treatment effects and survival, therefore, results in identifying more clinically relevant subgroups.
For a fair comparison, we present three alternatives of CMHE: one with fixed $K=3$ survival subgroups, one with $L=3$ treatment effect subgroups, and one with $K=L=2$, which allows for a total of 4 subgroups. Crucially, these methods assume proportional hazards for each subgroup and do not account for the treatment non-randomisation. Additionally, we also compare our model against its unadjusted alternative (CSC Unadjusted), which uses the unweighted factual likelihood ($w_{i}=1$) and should suffer from non-randomisation of treatment assignment.
Finally, we compare against two step-wise approaches in which clustering and treatment effect are modelled separately. First, we use an unsupervised clustering algorithm on the covariates, followed by a non-parametric estimate of the treatment effect as proposed in Nagpal et al. (2022a), referred as Kmeans + TE in the following. Specifically, we use a K-Means (Hartigan & Wong, 1979) to cluster the data and we compute the difference between Kaplan-Meier (Kaplan & Meier, 1958) estimates between control and treated patients stratified by clusters. However, this approach is limited by its implicit assumption of randomisation and its pre-specified number of subgroups, which is not informed by observed outcomes. Second, we use a Virtual Twins approach to estimate individualised treatment effects and then cluster them. As proposed in the literature, we propose to use a survival tree to estimate response under each treatment regime and then use a KMeans on the estimated treatment effects, computed as the difference between survival estimates. Note that this methodology also assumes treatment randomisation, and the clustering is uninformed by the covariates. For details on training and hyperparameter optimisation, we refer to App. B.2.
@@
In the synthetic experiments, the subgroup structure is known. We measure the adjusted [^6] Random patient assignment results in an adjusted Rand-Index of 0. Rand-Index (Rand, 1971), which quantifies how the estimated assignment aligns with the known underlying group structure. Additionally, we use the integrated absolute error (IAE) between the treatment effect estimate and the ground truth, which measures how well we recover each subgroup’s treatment effect
@@
where $\hat{\tau}_{\hat{k}}$ is the estimated treatment effect for subgroup $\hat{k}=\operatorname*{arg\,max}_{l}\mathbb{E}_{x\in k}(Z=l\mid x)$, i.e. the most likely assigned cluster for patients in the underlying $k$ cluster, and $\tau_{k}$ is the ground truth.
@@ 3.3 Treatment effect recovery
In this section, we present the results under the previously described data generation process. App. B.4 presents alternative datasets which demonstrate the flexibility of the proposed methodology.
@@
For all methodologies, we need to choose the number of subgroups a priori. An important question is, therefore, whether we can identify the true number of subgroups with our model. Figure 4 presents the average NLL obtained by cross-validation for models with different numbers of subgroups $K$. The dotted lines represent the elbow heuristic (Thorndike, 1953), which identifies a change point in the explained variability of a clustering strategy, here considering the log-likelihood. Using this heuristic, the optimal choice for $K$ is $3$, which agrees with the underlying generative process. App. B.3 explores the sensitivity to this choice, demonstrating the method’s robustness. This data-driven choice of the number of subgroups $K$ is a crucial strength of our method, compared to classical two-stage analyses which separate clustering from treatment effect estimates. In these methods, survival outcomes cannot directly guide the choice of $K$.
@@
Next, we present quantitative results of our method against the benchmarks methods. Table 1 presents the performance of the different methodologies under the two studied scenarios. Recall that $K$ denotes the number of treatment response subgroups, while in CMHE, the additional parameter $L$ describes the number of survival distributions.
Model Rand-Index $\text{IAE}_{k}(t_{max})$ $k=1$ $k=2$ $k=3$ Randomised CSC 0.798 (0.054) 0.042 (0.012) 0.032 (0.007) 0.015 (0.006) CSC Unadjusted 0.804 (0.052) 0.032 (0.009) 0.027 (0.007) 0.018 (0.010) CMHE ($K=3$) 0.392 (0.034) 0.164 (0.009) 0.077 (0.004) 0.070 (0.005) CMHE ($L=3$) 0.258 (0.155) - 0.088 (0.007) - CMHE ($K=L$) 0.392 (0.054) 0.338 (0.262) - - KMeans + TE 0.000 (0.002) 0.210 (0.011) 0.091 (0.007) 0.142 (0.016) Virtual Twins 0.578 (0.081) 0.022 (0.006) 0.032 (0.011) 0.025 (0.010) Observational CSC 0.812 (0.038) 0.037 (0.013) 0.034 (0.009) 0.022 (0.005) CSC Unadjusted 0.727 (0.084) 0.048 (0.019) 0.025 (0.011) 0.025 (0.009) CMHE ($K=3$) 0.385 (0.022) 0.169 (0.012) 0.078 (0.005) 0.075 (0.008) CMHE ($L=3$) 0.190 (0.127) 0.192 (0.010) - 0.140 (0.009) CMHE ($K=L$) 0.454 (0.068) 0.210 (0.013) 0.095 (0.020) 0.188 (0.014) KMeans + TE 0.001 (0.004) 0.192 (0.016) 0.090 (0.019) 0.147 (0.016) Virtual Twins 0.438 (0.139) 0.050 (0.039) 0.034 (0.033) 0.051 (0.033)
@@
Table 1: 5-fold cross-validated performance averaged (with standard deviation in parenthesis) under the Randomised and Observational treatment simulation. Best performance per column and simulation scenario are marked in bold, second best in italic. ’-’ describes when the methodology diverges. Our proposed CSC method and its unadjusted variant best recovers the underlying treatment responses, with the adjusted approach presenting the best performance in the observational setting.
@@
CSC outperforms all CMHE alternative, the current state-of-the-art method. CMHE’s parametrisation and assumptions explain this difference. Critically, CMHE assumes (i) proportional hazards and (ii) a linear impact of treatment on log-hazards. Neither of these two assumptions is likely to hold in real-world settings as mimicked by our synthetic data. In contrast, our approach does not constrain the treatment effect due to its flexible modelling of the survival function under both treatment regimes. Increasing the number of subgroups, as shown in $K=L$ improves the performance of CMHE in terms of clustering quality (Rand-Index) and the recovery of the underlying treatment effect (lower IAE), but still is inferior compared to our proposed methods. These experiments highlight the advantage of the proposed CSC method in uncovering subgroups of treatment responses due to the flexibility in modelling complex survival distributions under both treatment regimes without proportionality assumption.
CSC presents the best performance in identifying the underlying subgroup. Our proposed CSC method outperforms all approaches on the Rand-Index. In particular, the two-step approaches do not identify the underlying subgroups when covariates and outcomes subgroups are aligned. KMeans + TE identify subgroups that are independent of the outcome of interest, as shown by the Rand-Index. Consequently, the identified clusters differ in their treatment responses as shown by the large IAE. The Virtual Twins approach presents a better capacity to identify the clusters of interest with better Rand-Index and IAE. However, the two-step approach hurts subgroup identification due to the disconnect between the clustering and input covariates. This disconnect leads to a significantly lower Rand-Index in comparison to CSC despite accurate treatment effect estimates.
Treatment assignment correction improves subgroup identification in observational settings. The Observational simulation demonstrates the importance of correcting the likelihood under treatment non-randomisation. The two CSC alternatives present comparable performance in the randomised setting, as theoretically expected, due to a constant $w_{i}$ in this context. Note that estimating propensity slightly reduces performance due to error introduced, reducing performance for the corrected alternative. In a randomised setting, CSC Unadjusted should be preferred. However, the superiority of CSC over its unadjusted alternative is evident in the observational setting, where CSC better recovers the different groups as shown by the Rand-Index and subgroup treatment effects. Finally, Appendices B.4.1 to B.4.5 further validate the robustness of our method in settings with unequally distributed cluster sizes, population size, increased number of underlying clusters, different treatment rates and covariate structures.
Real-world heterogeneity analysis in adjuvant radiotherapy responses (App. C).
Using the Surveillance, Epidemiology, and End Results program [^7] Available at https://seer.cancer.gov/(Seer), App. C investigates a real-world setting: we analyse which subgroups of breast cancer patients most benefit from adjuvant radiotherapy, an active area of research. Our analysis highlights a subgroup characterised by a higher number of distant lymph nodes that benefit from treatment. Despite the inherent limitations of the dataset, our exploratory method identifies a subgroup that alternative methods miss, potentially informing future clinical trials to validate or challenge this hypothesis. Critically, this demonstrates that our work offers a practical methodology for exploratory subgroup analysis of medical data to identify treatment effect heterogeneity in observational studies.
@@ 4 Related work
While survival subgrouping has been proposed through mixtures of distributions (Nagpal et al., 2021a; b; Jeanselme et al., 2022) to identify different phenotypes of patients, the machine learning literature on phenotyping treatment effects with time-to-event outcomes remains sparse. The current literature focuses on estimating population or individual treatment effects (Curth et al., 2021; Shalit et al., 2017; Johansson et al., 2016; Zhang et al., 2020). While these models help to understand the average population-wide response to treatment and aim to estimate individualised treatment effects, they do not provide an understanding of the patient groups that may benefit more or be harmed more by treatment. Identifying such groups aligns with, and is thus more useful for, drafting medical guidelines that direct treatment to those subgroups most likely to benefit from it.
@@
Discovering intervenable subgroups is core to medical practice, particularly identifying subgroups of treatment effect, as patients do not respond like the average (Bica et al., 2021; Ruberg et al., 2010; Sanchez et al., 2022) and the average may conceal differential treatment responses. Identification of subgroups has long been used to design RCTs. Indeed, subgroups identified a priori can then be tested through trials (Cook et al., 2004; Rothwell, 2005). A posteriori analyses have gained traction to uncover subgroups of patients from existing RCTs to understand the underlying variability of responses.
@@
The first set of subgroup analysis methodologies consists of a step-wise approach: (i) estimate the ITE and (ii) uncover subgroups using a second model to explain the heterogeneity in ITE. Foster et al. (2011); Qi et al. (2021) describe the virtual twins approach in which one models the outcome using a decision tree for each treatment group. The difference between these decision trees results in the estimated treatment effects. A final decision tree aims to explain these estimated treatment effects to uncover subgroups. Similar approaches have been explored with different meta-learners (Xu et al., 2023), or Bayesian additive models (Hu et al., 2021), or replacing the final step with a linear predictor to uncover the feature influencing heterogeneity (Chernozhukov et al., 2018). However, Guelman et al. (2015) discuss the drawbacks associated with these approaches. Notably, the two-step optimisation may not lead to recovery of the underlying subgroups of treatment effects.
@@
Tree-based approaches were proposed to address the limitations of step-wise approaches by jointly discovering subgroups and modelling the treatment effect. Instead of traditional splits on the observed outcomes, these causal trees aim to discover homogeneous splits regarding covariates and treatment effects. Su et al. (2009) introduce a recursive population splitting based on the average difference in treatment effect between splits. Athey & Imbens (2015; 2016) improve the confidence interval estimation through the honest splitting criterion, which dissociates the splitting from the treatment effect estimation. Wager & Athey (2018) agglomerate these causal trees into causal forests for improved ITE estimation. Each obtained split in the decision tree delineates two subgroups of treatment effect Lipkovich et al. (2011); Loh et al. (2015). Alternatively, McFowland III et al. (2018) propose pattern detection and Wang & Rudin (2022), causal rule set learning to uncover these subgroups. However, all these approaches rely on a local optimisation criterion (Lipkovich & Dmitrienko, 2014) and greedy split exploration. Recently, Nagpal et al. (2020) addressed the local optimisation by constraining the treatment response to a linear form in a mixture of Cox models.
@@
Previous approaches uncover subgroups of treatment effect but consider RCTs with binary outcomes, not the observational setting with inherently continuous survival outcomes that our paper explores. At the intersection with survival analysis, Zhang et al. (2017) extend causal trees to survival causal trees, modifying the splitting criterion by measuring the difference in survival estimates between resulting leaves. Similarly, Hu et al. (2021) propose Bayesian additive models and Zhu & Gallego (2020) propose a step-wise approach with propensity weighting to study observational data. Closest to our work, (Jia et al., 2021; Nagpal et al., 2023) propose to uncover subgroups within RCTs with survival outcomes. Jia et al. (2021) propose a mixture of treatment effects characterised by Weibull distributions trained in an expectation-maximisation framework. Similarly, Nagpal et al. (2023) stratify the population into three groups: non-, positive- and negative responders to treatment. An iterative Monte Carlo optimisation is used to uncover these subgroups, characterised by a Cox model with a multiplicative treatment effect. As demonstrated in our work, this step-wise optimisation may be limiting, and the assumption of RCTs renders the model less relevant in observational data.
@@ 5 Conclusion
This work fills the current gap in the literature of methodologies to identify subgroups of treatment responses in observational time-to-event data. While modelling the observed outcome, the approach jointly identifies patient subgroups with different treatment responses. Our experiments demonstrate the capacity of our proposed method to uncover groups of treatment effects in both synthetic and real-world datasets. However, causal modelling relies on empirically unverifiable assumptions. The invalidity of these assumptions is a lesser concern in our targeted setting of hypothesis generation, in comparison to individual estimation of treatment recommendations.
Serving the purpose of a hypothesis-generating tool, we invite practitioners to further investigate the subgroups identified in observational studies through RCTs to (i) validate the estimated responses, (ii) identify potential alternative treatments with improved responses, and (iii) inform clinical guidelines.
Acknowledgements
The authors acknowledge the partial support of the UKRI Medical Research Council (MC_UU_00002/5 and MC_UU_00002/2). VJ, CHY and FF acknowledge the Enrichment Scheme of The Alan Turing Institute under the EPSRC Grant EP/N510129/1. FF acknowledges the receipt of studentship awards from the Health Data Research UK-The Alan Turing Institute Wellcome PhD Programme (Grant Ref: 218529/Z/19/Z).